Assume that the paired data came from a population that is normally distributed. Using a

0.05


significance level and

d=x−​y,

find

_
d,

sd​,

the t test​ statistic, and the critical values to test the claim that

μd=0.

x


13


7


8


6


6


14


8


9


 

y


13


7


5


5


6


9


9


12


_
d=

 

​(Round to three decimal places as​ needed.)

sd=

 

​(Round to three decimal places as​ needed.)

t=

 

​(Round to three decimal places as​ needed.)

tα/2=±

 

​(Round to three decimal places as​ needed.)
Answer #1

SOLUTION:

From given data,

Test hypothesis

Null hypothesis:

Ho :μd=0

Alternative Hypothesis

Ha: μd≠0

From given values,

X Y

d=

X-Y

\bar{d} d-\bar{d} (d-\bar{d})2
13 13 13-13=0 0.625 0-0.625 = -0.625

(-0.625)2

= 0.390625

7 7 7-7 = 0 0.625 0-0.625 = -0.625

(-0.625)2

= 0.390625

8 5 8-5 = 3 0.625 3-0.625 = 2.375

(2.375)2

= 5.640625

6 5 6-5 = 1 0.625 1-0.625 = 0.375

( 0.375)2

= 0.140625

6 6 6-6=0 0.625 0-0.625 = -0.625

(-0.625)2

= 0.390625

14 9 14-9= 5 0.625 5-0.625 = 4.375

(4.375)2

= 19.140625

8 9 8-9=-1 0.625 -1-0.625 =-1.625

(-1.625)2

= 0.390625

9 12 9-12 = -3 0.625 -3-0.625 = -3.625

(-3.625)2

= 13.140625

\Sigma d = 5

\Sigma (d-\bar{d})2

= 39.234375

\bar{d} = \Sigma d / n = 5/8 = 0.625

Standard deviation.

Compute the standard deviation (sd) of the differences computed from n matched pairs.

sd = sqrt [ (Σ(d - \bar{d})2 / (n - 1) ]

sd = sqrt [ 39.234375/ (8 - 1) ]

sd = sqrt [ 39.234375/ 7 ]

sd = sqrt [ 35.6049107]

sd = 2.367

Test statistic.

The test statistic is a t statistic (t) defined by the following equation.

t = (\bar{d} - \mu) / SE

the standard error can be approximated by:

SE = sd / sqrt( n )

=2.367/sqrt(8)

=0.83686

t = (0.625 - 0) / 0.83686

t = 0.625 / 0.83686

t = 0.7468

Critical value :

Degrees of freedom.

The degrees of freedom (DF) is: DF = n - 1 =8-1=7

significance level = \alpha = 0.05

t\alpha/2 ,df = t0.05/2 ,7 = t0.025 ,7 = 2.365

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.